NERSCPowering Scientific Discovery Since 1974

January 4

PDSF users meeting 1/4/11

Attending: Eric, Katie and Jay from NERSC and users Jeff P., Thomas, Art and possibly an an unknown person on the phone (?).

Cluster Status/Utilization:  Quite full over the break and stable.  Only issue was that at one point there were 26k jobs running or pending but that didn't cause any major problems.

Outages:  None.

Upcoming downtimes:  Need to do some work on rack 22 - clean up and rearrange some things.  This will affect STAR db and will take ~5 hours.  eliza16/17/18 will also get worked on.  HPSS downtime 1/12.  Also will need to restart SGE at some point.

Procurements/New Hardware:  eliza5 additional storage is being worked on - had to check with vendor to be sure the new hardware would work with the existing hardware.  Requisition will go out soon.

sl302 retirement: still waiting for a downtime when the interactives can be rebooted.

Topics from users:

- Some discussion about chos documentation

- Discussed a bit about how to manage SRM.

The rest of the meeting was a demo of the new website by Eric.